# Approximation in Hölder spaces

### v3

@@ Appendix A Comparison of $\dot{C}^{0,\omega}$ and uniform convergence
Indeed, let $B$ a bounded subset of $X.$ Given $\varepsilon>0,$ let $N$ be so that $|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon/(1+\omega(\operatorname{diam}(B)))$ for all $n\geq N.$ Then, for those $n\geq N,$ and $x\in B,$ we can write
@@
$\leq\left(1+\omega(\|x\|)\right)|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\left(1+\omega(\operatorname{diam}(B))\right)|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon.$
@@
Let $\omega$ be of the form $\omega(t)=t^{\alpha},$ with $0<\alpha\leq 1$ Define $f_{n}:\mathbb{R}\to\mathbb{R}$ as the continuous function given by $f_{n}=0$ on $(-\infty,0]\cup[2n,+\infty),$ $f_{n}(n)=1$, and $f_{n}$ affine on both $[0,n]$ and $[n,2n].$ It is immediate that $\|f_{n}\|_{\infty}\leq 1$ for all $n,$ and that $(f_{n})_{n}$ does not converge to 00 uniformly on $\mathbb{R}.$ However, one has $f_{n}\to 0$ in the $\dot{C}^{0,\omega}(\mathbb{R})$ sense. This easily follows after observing that
@@
Finally, we next show that the converse to Remark A.1 is false, by providing a sequence $(f_{n})_{n}\subset\dot{C}^{0,\omega}$ and $f\in\dot{C}^{0,\omega}$ so that $f_{n}$ converges uniformly on $\mathbb{R},$ and $f_{n}$ does not converge to $f$ in the $\|\cdot\|_{\dot{C}^{0,\omega}}$ norm on bouded subsets.

### v4

@@ Appendix A Comparison of $\dot{C}^{0,\omega}$ and uniform convergence
Indeed, let $B$ a bounded subset of $X$ with $0\in B.$ Given $\varepsilon>0,$ let $N$ be so that $|f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon/(1+\omega(\operatorname{diam}(B)))$ for all $n\geq N.$ Then, for those $n\geq N,$ and $x\in B,$ we can write
@@
[…]
@@
Let $\omega$ be of the form $\omega(t)=t^{\alpha},$ with $0<\alpha\leq 1.$ Define $f_{n}:\mathbb{R}\to\mathbb{R}$ as the continuous function given by $f_{n}=0$ on $(-\infty,0]\cup[2n,+\infty),$ $f_{n}(n)=1$, and $f_{n}$ affine on both $[0,n]$ and $[n,2n].$ It is immediate that $\|f_{n}\|_{\infty}\leq 1$ for all $n,$ and that $(f_{n})_{n}$ does not converge to 0 uniformly on $\mathbb{R}.$ However, one has $f_{n}\to 0$ in the $\dot{C}^{0,\omega}(\mathbb{R})$ sense. This easily follows after observing that
@@
Finally, we next show that the converse to Remark A.1 is false, by providing a sequence $(f_{n})_{n}\subset\dot{C}^{0,\omega}$ and $f\in\dot{C}^{0,\omega}$ so that $f_{n}$ converges uniformly on $\mathbb{R},$ and $f_{n}$ does not converge to $f$ in the $\|\cdot\|_{\dot{C}^{0,\omega}}$ norm on bounded subsets.